# Given Enough Time

### v1

@@ Abstract
Deep learning is closing the gap with humans on several object recognition benchmarks. Here we investigate this gap in the context of challenging images where objects are seen from unusual viewpoints. We find that humans excel at recognizing objects in unusual poses, in contrast with state-of-the-art pretrained networks (EfficientNet, SWAG, ViT, SWIN, BEiT, ConvNext) which are systematically brittle in this condition. Remarkably, as we limit image exposure time, human performance degrades to the level of deep networks, suggesting that additional mental processes (requiring additional time) take place when humans identify objects in unusual poses. Finally, our analysis of error patterns of humans vs. networks reveals that even time-limited humans are dissimilar to feed-forward deep networks. We conclude that more work is needed to bring computer vision systems to the level of robustness of the human visual system. Understanding the nature of the mental processes taking place during extra viewing time may be key to attain such robustness.
@@ 1 Introduction
In an era marked by the rapid advancement of deep learning in computer vision, a natural question arises: Can machines meet and even exceed the capabilities of the human visual system? Numerous recent studies have shown that deep networks outperform humans on well-known object recognition benchmarks (e.g., ImageNet: He et al. (2015); Vasudevan et al. (2022); Dehghani et al. (2023)). Deep networks are even said to outperform humans on out-of-distribution recognition tasks (Geirhos et al., 2021), where the images used for evaluation are subjected to distortions that the networks were not specifically trained on.
However, studies that investigate out-of-distribution generalization mostly focus on local distortions (e.g., texture modifications, blur, color modifications) as opposed to transformations that affect the global structure of the image, such as a change of viewpoint. A few studies have investigated the generalization capability of deep networks to recognize objects in various poses (Alcorn et al., 2019; Ibrahim et al., 2022; Abbas & Deny, 2023), and they show a substantial degradation of network performance for unusual poses. However, a human benchmark has been missing from these studies, leaving the question open of whether humans outperform networks on this task.
Here, we compare human subjects with state-of-the-art deep networks at recognizing objects in various poses. Remarkably, we show that—while both humans and networks excel on upright poses—humans substantially outperform networks on unusual poses. However, as we limit viewing time, humans exhibit a similar brittleness to deep networks on unusual poses, demonstrating the necessity of additional mental processes for analysing these challenging images. Although networks and time-limited humans are both brittle, an analysis of their patterns of error reveals that time-limited humans make different mistakes than networks. Altogether, our results show that (1) humans are still much more robust than networks at recognizing objects in unusual poses, (2) time is of the essence for such ability to emerge, and (3) even time-limited humans are dissimilar to deep neural networks.
@@ 2.1 Dataset collection
We chose 51 different object categories from the ImageNet classes (see Appendix B for the list of objects), and obtained a corresponding 3D model for each category from Sketchfab (https://sketchfab.com/). The objects were chosen based on a few criteria:
@@
  • The object in its upright pose should be correctly labeled by Noisy Student EfficientNet-L2, pretrained on JFT-300M and finetuned on ImageNet (Xie et al., 2020) (referred to as EfficientNet below). We chose EfficientNet to guide our selection of objects because it is the best open-source network at recognizing objects in unusual poses, according to a recent comprehensive study comparing 43 state-of-the-art networks on this task (Abbas & Deny, 2023).
@@ 2.2.1 Observers
Altogether 24 observers (12 women, 12 men, aged 19-54) participated in the experiments. Participants were required to have normal vision or corrected-to-normal vision, and be fluent in Finnish. Additionally, the observers did not report any medical condition (epilepsy, migraine, etc.) that could affect the results.
@@ 2.2.3 Procedure
Each observer participated in a limited viewing time experiment and an unlimited viewing time experiment. In the limited viewing time experiment, each trial proceeded as follows (see Figure 2a). First, a fixation cross-hair was presented in the centre of the screen. Once the observer was ready, they pressed the space bar. After 500 ms, the object image was presented. The duration of the object image presentation was 40 ms for 12 observers, and 150 ms for the other 12. These stimulus durations (40 and 150 ms) were based on pilot experiments which suggested that those durations would lead to performance levels between perfect and chance level performance. The object image was immediately followed by 500 ms of dynamic chromatic white noise. After that, two response alternatives (in Finnish), one of them always the correct one, were presented and the observer had to choose (by pressing one of two keys on the keyboard) which alternative corresponded to the presented object. Each observer performed 49 trials, in which the image was in one of three types of poses: upright in 17 trials, rotated-correct (correctly classified by EfficientNet, see Dataset collection 2.1) in 17 trials, and rotated-incorrect (incorrectly classified by EfficientNet) in 15 trials. The trial types were interleaved and presented in random order. The unlimited viewing time experiment (see Figure 2b) followed the same structure, except that the image remained on the screen until the observer responded (2.0 s on average), the labels were present from the start, and no noise was presented after the object image.
@@ 2.3 Machine tests
We implemented the same tests on a collection of 5 state-of-the-art deep networks. Since EfficientNet was used to select the problematic images and labels, we did not include EfficientNet in this collection nor in our study of network robustness. We, in fact, investigated the transfer of errors between EfficientNet and these other networks.
We tested the 5 following deep networks: SWAG-RegNetY (Singh et al., 2022), BEiT-L/16 (Bao et al., 2021), ConvNext-XL (Liu et al., 2022), SWIN-L (Liu et al., 2021), and ViT-L/16 (Dosovitskiy et al., 2020) (see Appendix A for model descriptions). The network selection was based on previous work from Abbas & Deny (2023), which thoroughly evaluated the robustness of 43 state-of-the-art networks to objects presented in unusual poses. This collection provides a diverse set of state-of-the-art and very large models trained on large datasets (e.g., Imagenet-21K, Instagram 3.6B) with various architectures (ConvNet and Transformers) and training objectives (supervised and self-supervised). These networks obtained some of the most robust results to unusual poses according to Abbas & Deny (2023), and all of them were fine-tuned to the categories of ImageNet.
We mirrored the experiment done on humans by submitting each network to the same two-forced-choice test. Each network was shown the 147 different images of objects, 51 upright poses, 51 rotated-correct poses, and 45 rotated-incorrect poses. We compared the activations of the units of the final layer corresponding to each of the two categories given as choices for each image, and selected the category with corresponding highest unit-activation among the two to be the chosen answer.
@@ 2.4 Statistical tests
Figure 3: Comparing neural networks and humans at recognizing objects in various poses. Grey bars: average deep network performance (left: upright vs. right: rotated) with 95% confidence intervals (n=5). Light grey points are individual networks. Chance performance is 50%. Blue bars: human averages with unlimited viewing time (n=12). Orange & Red bars: human averages with limited viewing time (150 ms and 40 ms respectively, n=12). Three stars indicate highly significant differences (p$<$0.001), ”n.s.” for not significant. With unlimited time, humans excel at recognizing rotated objects, while networks struggle (best: SWAG, 70.1%). Limiting viewing time (40 ms or 150 ms) impairs human performance for rotated objects, substantially more than for upright objects.
@@ 3 Results
We compared human and machine vision on a task consisting in recognizing objects in unusual poses. For this, we selected 5 different deep neural networks (SWAG, ViT, SWIN, BEiT, ConvNext), chosen for their state-of-the-art performance in recognizing objects in unusual poses (Abbas & Deny, 2023). We performed a two-forced-choice object categorization task: an image of an object was presented to the viewer (either deep network or human participant), and then the viewer had to choose between two different names (i.e., labels) for the object. For networks, the choice was made by looking at the highest activation of the softmax output layer for these two labels. The two labels were selected based on the highest activations of Noisy Student EfficientNet’s output layer (Xie et al., 2020), the most robust network to unusual poses according to Abbas & Deny (2023). EfficientNet was excluded from our analyses comparing humans and networks because images and labels were chosen based on its mistakes (see Methods).
@@
With unlimited viewing time, humans outperform neural networks at recognizing objects in unusual poses.
@@
First, we compared the ability of humans and networks to recognize objects in upright and unusual poses when humans were not given any time limit for the object recognition (Figure 3). In the upright poses, both humans and neural networks performed very well: 4 networks obtained 100% accuracy and one (ViT-L) failed to correctly label only 1 out of 51 objects; humans had slightly lower performance with an accuracy of 98.0% $\pm$ 2.4% (95% confidence intervals). The very few errors that human observers made are likely to be keystroke errors and misunderstandings of the verbal response alternatives (as reported by some subjects). For unusual poses, human participants with unlimited viewing time clearly outperformed neural networks. For humans, the performance on rotated images and upright images was on par (rotated condition: 98.9% $\pm$ 1.0%)(t(11) = -0.79, p = 0.45), whereas for networks, the performance drastically dropped between upright (99.6% $\pm$ 1.0%) and rotated objects (67.1% $\pm$ 4.3%), with a drop of over 30% performance on average (t(4) = 18.7, p = 4.8e-05). Even the most robust network of our collection, SWAG trained on IG-3.6B (Singh et al., 2022), showed a drop in performance of 30% (from 100% on upright poses to 70.1% on rotated poses).
@@
Figure 4: Error patterns are different for neural networks and time-limited humans. a) Comparison of human and network accuracy for the rotated-correct condition (EfficientNet was correct on these rotations) vs. rotated-incorrect condition (rotations that have failed EfficientNet). EfficientNet errors transfer well to other networks, but not to humans. b) Error consistency analysis performed on the 5 neural networks and 40 ms time-limited human subjects. 12 human subjects were partitioned into 4 groups of 3, so that every group saw each rotated image exactly once. The darker red cluster for networks (dark red = highly consistent errors) indicates that they have similar patterns of error, which are not shared by human subjects. c) Mean error consistencies were calculated by comparing networks with each other and with human subjects (i.e., average computed over the different matrix clusters). A two-tailed unpaired t-test confirms that networks indeed make more consistent errors with each other than with humans (t(28) = 4.0, p = 3.8e-04).
@@
Figure 5: Examples of objects in unusual poses, where networks and 40 ms time-limited humans made similar and differing errors. Above each image is a quantitative score of correct answers and below each image are the given answer choices, correct answer being the first. a) Images where all networks and humans correctly labeled the object. b) Images where all networks failed to correctly label the objects, but where humans mostly chose the correct answer. c) Images, where all the networks correctly labeled the objects, but where most humans failed. d) Images, where both networks and humans were mostly not able to correctly label the object.
@@
Patterns of errors differ between humans and networks.
@@
We then sought to understand what qualitatively explained the differences in the error patterns between time-limited humans (40 ms) and networks.

### v2

# A comparison between humans and AI at recognizing objects in unusual poses

@@ Abstract
Deep learning is closing the gap with human vision on several object recognition benchmarks. Here we investigate this gap for challenging images where objects are seen in unusual poses. We find that humans excel at recognizing objects in such poses. In contrast, state-of-the-art deep networks for vision (EfficientNet, SWAG, ViT, SWIN, BEiT, ConvNext) and state-of-the-art large vision-language models (Claude 3.5, Gemini 1.5, GPT-4) are systematically brittle on unusual poses, with the exception of Gemini showing excellent robustness in that condition. As we limit image exposure time, human performance degrades to the level of deep networks, suggesting that additional mental processes (requiring additional time) are necessary to identify objects in unusual poses. An analysis of error patterns of humans vs. networks reveals that even time-limited humans are dissimilar to feed-forward deep networks. In conclusion, our comparison reveals that humans and deep networks rely on different mechanisms for recognizing objects in unusual poses. Understanding the nature of the mental processes taking place during extra viewing time may be key to reproduce the robustness of human vision in silico.
@@ 1 Introduction
In an era marked by the rapid advancement of deep learning for computer vision, a natural question arises: Can machines meet or even exceed the capabilities of the human visual system? Numerous recent studies have shown that deep networks outperform humans on well-known object recognition benchmarks (e.g., ImageNet: He et al. (2015); Vasudevan et al. (2022); Dehghani et al. (2023)). Deep networks are even said to outperform humans on out-of-distribution recognition tasks (Geirhos et al., 2021), where the images used for evaluation are subjected to distortions that the networks were not specifically trained on.
However, studies that investigate out-of-distribution generalization mostly focus on local distortions (e.g., texture modifications, blur, color modifications) as opposed to transformations that affect the global structure of the image, such as a change in viewpoint. A few studies have investigated the generalization capability of deep networks to recognize objects in various poses, both in non-adversarial (Alcorn et al., 2019; Ibrahim et al., 2022; Abbas & Deny, 2023; Madan et al., 2022) and adversarial settings (Zeng et al., 2019; Madan et al., 2021), and they show a substantial degradation of network performance for unusual poses. However, a direct comparison with human vision has been missing from these studies, leaving open the question of whether humans outperform networks on this task.
Here, we compare human subjects with state-of-the-art deep networks for vision and state-of-the-art large vision-language models at recognizing objects in various poses. We show that—while both humans and networks excel on upright poses—humans substantially outperform networks on unusual poses (except for Google’s Gemini showing exceptional robustness in that condition). As we limit viewing time, humans exhibit a similar brittleness to deep networks on unusual poses, demonstrating the necessity of additional mental processes for analysing these challenging images. Although networks and time-limited humans are both brittle, an analysis of their patterns of error reveals that time-limited humans make different mistakes than networks. In conclusion, our results show that (1) humans are still much more robust than most networks at recognizing objects in unusual poses, (2) time is of the essence for such ability to emerge, and (3) even time-limited humans are dissimilar to deep neural networks.
@@ 2.1 Dataset collection
We chose 51 different object categories from the ImageNet classes (see Appendix C for the list of objects), and obtained a corresponding 3D model for each category from Sketchfab (https://sketchfab.com/). The objects were chosen based on a few criteria:
@@
  • The object in its upright pose should be correctly labeled by Noisy Student EfficientNet-L2, pretrained on JFT-300M and finetuned on ImageNet (Xie et al., 2020) (referred to as EfficientNet below). We chose EfficientNet to guide our selection of objects because it is the best open-source network at recognizing objects in unusual poses, according to a recent comprehensive study comparing 37 state-of-the-art networks on this task (Abbas & Deny, 2023).
@@
The stimulus set applied here is not entirely ecologically realistic, as all objects are displayed without a background. However, our aim was to create a highly controlled stimulus set and experimental setup. Performance in canonical view is compared with performance with exactly the same object in a rotated view, both shown without background. Thus, any performance differences, be it human or network, can only be caused by rotation. Conversely, any robustness can only be caused by rotation-invariance, not context-based reasoning.
@@ 2.2.1 Observers
Altogether 24 observers (12 women, 12 men, aged 19-54) participated in the experiments. Our observers comprised Finnish students from diverse academic backgrounds, including disciplines such as engineering, psychology, business, and medicine. Participants were required to have normal vision or corrected-to-normal vision. Additionally, the observers did not report any medical condition (epilepsy, migraine, etc.) that could affect the results.
@@ 2.2.3 Procedure
Prior to the experiments, observers were provided with a small training set, in order to familiarize themselves with the experimental procedure. None of the training objects were featured in the real experiment. After this training, each observer participated in a limited viewing time experiment and an unlimited viewing time experiment. In the limited viewing time experiment, each trial proceeded as follows (see Figure 2a). First, a fixation cross-hair was presented in the centre of the screen. Once the observer was ready, they pressed the space bar. After 500 ms, the object image was presented. The duration of the object image presentation was 40 ms for 12 observers, and 150 ms for the other 12. These stimulus durations (40 and 150 ms) were based on pilot experiments which suggested that those durations would lead to performance levels between perfect and chance level performance. The object image was immediately followed by 500 ms of dynamic chromatic white noise. After that, two response alternatives (in Finnish), one of them always the correct one, were presented and the observer had to choose (by pressing one of two keys on the keyboard) which alternative corresponded to the presented object. Each observer performed 49 trials, in which the image was in one of three types of poses: upright in 17 trials, rotated-correct (correctly classified by EfficientNet, see Dataset collection 2.1) in 17 trials, and rotated-incorrect (incorrectly classified by EfficientNet) in 15 trials. The trial types were interleaved and presented in random order. The unlimited viewing time experiment (see Figure 2b) followed the same structure, except that the image remained on the screen until the observer responded (2.0 s on average), the labels were present from the start, and no noise was presented after the object image.
@@ 2.3 Machine tests
We implemented the same tests on a collection of 5 state-of-the-art deep networks for vision (referred below as ’pure vision networks’) and 6 large vision-language models (VLM). Since EfficientNet was used to select the problematic images and labels, we did not include EfficientNet in this collection nor in our study of network robustness. We, in fact, investigated the transfer of errors between EfficientNet and these other networks.
We tested the 5 following pure vision networks: SWAG-RegNetY (Singh et al., 2022), BEiT-L/16 (Bao et al., 2021), ConvNext-XL (Liu et al., 2022), SWIN-L (Liu et al., 2021), and ViT-L/16 (Dosovitskiy et al., 2020) (see Appendix B for model descriptions). The network selection was based on previous work from Abbas & Deny (2023), which thoroughly evaluated the robustness of 37 state-of-the-art networks to objects presented in unusual poses. This collection provides a diverse set of state-of-the-art and very large models trained on large datasets (e.g., Imagenet-21K, Instagram 3.6B) with various architectures (ConvNet and Transformers) and training objectives (supervised and self-supervised). These networks obtained some of the most robust results to unusual poses according to Abbas & Deny (2023), and all of them were fine-tuned to the categories of ImageNet.
We mirrored the experiment done on humans by submitting each pure vision network to the same two-forced-choice test. Each network was shown the 147 different images of objects, 51 upright poses, 51 rotated-correct poses, and 45 rotated-incorrect poses. We compared the activations of the units of the final layer corresponding to each of the two categories given as choices for each image, and selected the category with corresponding highest unit-activation among the two to be the chosen answer.
Our selection of VLMs included Gemini 1.5 Flash, Gemini 1.5 Pro, Claude 3 Opus, Claude 3.5 Sonnet, GPT-4o and GPT-4-vision-preview. The experiment was conducted via the API. Each model was shown the 147 images and provided with the following prompt (see examples in Appendix A):
What’s in this image?
A. [label 1]
B. [label 2]
Choose either A or B and answer in one or two words.
If the model chose the correct label, the answer was counted as correct. If the model chose the wrong label or gave a response outside of the label options, the answer was counted as incorrect. If the model didn’t provide any answer (e.g., Gemini threw a safety error for a few images), the image was disregarded and not included in the analysis.
@@ 2.4 Statistical tests
Figure 3: Comparing neural networks and humans at recognizing objects in various poses. Dark grey bars: Average performance of pure vision deep networks (left: upright vs. right: rotated) with 95% confidence intervals (n=5, grey points represent individual network performances). Light grey bars: Average performance of six large vision-language models (VLMs). Diamonds indicate individual model performances (full diamonds show the best-performing version of the different model classes). Blue bars: Average human performance with unlimited viewing time (n=12, grey points represent individual performances). Orange & Red bars: Average human performance with limited viewing time (150 ms and 40 ms, respectively, n=12, grey points represent individual performances). Chance performance is 50%. Three stars (***) indicate highly significant differences (p$<$0.001), ”n.s.” for not significant. With unlimited time, humans excel at recognizing rotated objects, while pure vision networks struggle (best: SWAG at 70.1%). Both GPT-4 and Claude models follow the same pattern as the pure vision networks, with a significant drop in accuracy for rotated images compared to upright ones. However, Gemini 1.5 Flash mirrors human performance with unlimited viewing time, achieving 97.7% accuracy on rotated images compared to human accuracy of 98.9%. Limiting human viewing time (40 ms or 150 ms) impairs their ability to recognize rotated objects, substantially more than upright objects, bringing their performance closer to the network level.
@@ 3 Results
We compared human and machine vision on a task consisting in recognizing objects in unusual poses. For this, we selected 5 different pure vision deep neural networks (SWAG, ViT, SWIN, BEiT, ConvNext), chosen for their state-of-the-art performance in recognizing objects in unusual poses (Abbas & Deny, 2023), and 6 state-of-the-art large vision-language models (Gemini 1.5 Flash, Gemini 1.5 Pro, Claude 3 Opus, Claude 3.5 Sonnet, GPT-4o, GPT-4-vision-preview). We performed a two-alternative-forced-choice object categorization task: an image of an object was presented to the viewer (either deep network or human participant), and then the viewer had to choose between two different names (i.e., labels) for the object. The two labels were selected based on the highest activations of Noisy Student EfficientNet’s output layer (Xie et al., 2020), the most robust network to unusual poses according to Abbas & Deny (2023). For pure vision networks, the choice was made by looking at the highest activation of the softmax output layer for these two labels. EfficientNet was excluded from our analyses comparing humans and networks because images and labels were chosen based on its mistakes (see Methods). For the vision-language models, each VLM was given an image and prompted with the 2 label choices, and the label chosen by the model was selected to be its answer. The performance of VLMs is discussed in the last paragraph of results, while previous paragraphs concentrate on the comparison of pure vision networks and humans.
@@
With unlimited viewing time, humans outperform deep networks at recognizing objects in unusual poses.
@@
First, we compared the ability of humans and networks to recognize objects in upright and unusual poses when humans were not given any time limit for the object recognition (Figure 3). In the upright poses, both humans and neural networks performed very well: 4 networks obtained 100% accuracy and one (ViT-L) failed to correctly label only 1 out of 51 objects; humans had slightly lower performance with an accuracy of 98.0% $\pm$ 2.4% (95% confidence intervals). The very few errors that human observers made are likely to be keystroke errors and misunderstandings of the verbal response alternatives (as reported by some subjects). We conducted an additional test to discern whether network errors arise from changes in viewpoint, or alterations in image statistics between training and testing distributions (natural vs. synthetic). We subjected all five of our networks to a dataset comprising objects rotated at angles of ±$\pm 10^{\circ}$ from their canonical poses. All of our networks obtained perfect or near-perfect results (SWAG-RegNetY: 100%, BEiT-L/16: 99.3%, ConvNext-XL; 98.2%, SWIN-L: 91.1%, ViT-L/16: 95.0%). For unusual poses, human participants with unlimited viewing time clearly outperformed neural networks. For humans, the performance on rotated images and upright images was on par (rotated condition: 98.9% $\pm$ 1.0%)(t(11) = -0.79, p = 0.45), whereas for networks, the performance drastically dropped between upright (99.6% $\pm$ 1.0%) and rotated objects (67.1% $\pm$ 4.3%), with a drop of over 30% performance on average (t(4) = 18.7, p = 4.8e-05). Even the most robust network of our collection, SWAG trained on IG-3.6B (Singh et al., 2022), showed a drop in performance of 30% (from 100% on upright poses to 70.1% on rotated poses).
@@
Figure 4: Error patterns are different for neural networks and time-limited humans. a) Comparison of human and network accuracy for the rotated-correct condition (EfficientNet was correct on these rotations) vs. rotated-incorrect condition (rotations that have failed EfficientNet). Humans show consistent accuracy between rotated-correct and rotated-incorrect conditions. In contrast, networks, including the six VLMs (represented by diamonds), exhibit a performance drop. However, for Gemini 1.5 Flash, Gemini 1.5 Pro and Claude Opus, the decrease in performance between the two conditions is not as notable. b) Error consistency analysis performed on the 5 neural networks (not including VLMs) and 40 ms time-limited human subjects. 12 human subjects were partitioned into 4 groups of 3, so that every group saw each rotated image exactly once. The darker red cluster for networks (dark red = highly consistent errors) indicates that they have similar patterns of error, which are not shared by human subjects, highlighting that EfficientNet errors transfer better to other networks than to humans. c) Mean error consistencies were calculated by comparing networks with each other and with human subjects (i.e., average computed over the different matrix clusters). A two-tailed unpaired t-test confirms that networks indeed make more consistent errors with each other than with humans (t(28) = 4.0, p = 3.8e-04).
@@
Figure 5: Examples of objects in unusual poses, where deep networks for pure vision and 40 ms time-limited humans made similar and differing errors. Above each image is a quantitative score of correct answers and below each image are the given answer choices, correct answer being the first. a) Images where all networks and humans correctly labeled the object. b) Images where all networks failed to correctly label the objects, but where humans mostly chose the correct answer. c) Images, where all the networks correctly labeled the objects, but where most humans failed. d) Images, where both networks and humans were mostly not able to correctly label the object.
@@
Patterns of errors differ between humans and deep networks.
@@
We then sought to understand what qualitatively explained the differences in the error patterns between time-limited humans (40 ms) and deep networks.
@@
Next, we tested six large vision-language models (VLMs): GPT-4o, GPT-4-vision-preview, Claude 3.5 Sonnet, Claude 3 Opus, Gemini 1.5 Flash and Gemini 1.5 Pro.
With upright poses, all 6 models performed well (Gemini 1.5 Flash, GPT-4o and GPT-4-vision-preview achieving 100% accuracy, Gemini 1.5 Pro 98.0%, Claude 3.5 Sonnet 94.1% and Claude 3 Opus 92.2.%). However, there were significant differences in performance when recognizing rotated poses (Figure 3). Claude 3 Opus performed below the 50% chance level, achieving only 45.8% accuracy. Claude 3.5 Sonnet’s accuracy was on par with pure vision networks with 62.5% accuracy. GPT-4o had an accuracy of 84.8% and GPT-4-vision-preview 79.7%, performing substantially better than our best performing pure vision model, SWAG (70.1%). Remarkably, Gemini 1.5 Pro achieved an accuracy of 96.6% and Gemini 1.5 Flash an accuracy of 97.7.%, performances on par with unlimited humans (98.9%). Separating the rotated condition into two categories—rotated-correct (images correctly labelled by EfficientNet) and rotated-incorrect (images incorrectly labelled by EfficientNet)—revealed that the 6 VLMs exhibited a performance drop between these conditions, similarly to the five pure vision networks (Figure 4a). Performance differences of GPT-4o (18%), GPT-4-vision-preview (17%) and Claude 3.5 Sonnet (32%) were the most similar to the 5 pure vision networks, whose average performance gap was 26%. For Gemini models and Claude 3 Opus, the performance differences were much smaller: 4.7% for Gemini 1.5 Flash, 2.7% for Gemini 1.5 Pro and 2.4% for Claude. For both unlimited and time-limited humans, the performance difference was not significant between the two conditions. This suggests that errors from EfficentNet transfer better to most VLMs than to humans.